# The Hierarchical Growth of Bright Central Galaxies and Intracluster Light as Traced by the Magnitude Gap

### v1

@@ 7 Conclusion
Going forward, there remain many research questions we aim to explore to improve our observational understanding and characterisation of the BCG+ICL. In particular, understanding how the BCG+ICL stellar mass and radial distribution evolves with redshift. In future analysis using the Euclid Mission (as demonstrated by Kluge et al. (2024)) and the Rubin Observatory’s LSST, we plan to extend this type of analysis out to higher redshifts () to investigate the formation and evolution of the BCG+ICL system and how its correlation with halo mass evolves. Moreover, given the correlation between M14 and ICL fraction, we suggest that characterizing large M14 systems in Euclid and LSST, or targetting them with JWST, may provide information about the earliest appearances of the ICL. Additionally, we aim to determine whether the correlation between M14 and BCG stellar mass persists out to high redshift and whether this is intrinsic (i.e., is it in place prior to ex-situ growth) or results entirely from hierarchical formation.
@@ Acknowledgements
JBGM would like to thank Nina Hatch, the members of the NottICL research group, and the participants in the ISSI ICL working group for many useful discussions about the nature of ICL and measurement techniques. JBGM gratefully acknowledges support from the Leverhulme Trust.
@@
The Collaborating Institutions are Argonne National Laboratory, the University of California at Santa Cruz, the University of Cambridge, Centro de Investigaciones Energéticas, Medioambientales y Tecnológicas-Madrid, the University of Chicago, University College London, the DES-Brazil Consortium, the University of Edinburgh, the Eidgenössische Technische Hochschule (ETH) Zürich, Fermi National Accelerator Laboratory, the University of Illinois at Urbana-Champaign, the Institut de Ciències de l’Espai (IEEC/CSIC), the Institut de Física d’Altes Energies, Lawrence Berkeley National Laboratory, the Ludwig-Maximilians Universität München and the associated Excellence Cluster Universe, the University of Michigan, NSF’s NOIRLab, the University of Nottingham, The Ohio State University, the University of Pennsylvania, the University of Portsmouth, SLAC National Accelerator Laboratory, Stanford University, the University of Sussex, Texas A&M University, and the OzDES Membership Consortium.
Based in part on observations at Cerro Tololo Inter-American Observatory at NSF’s NOIRLab (NOIRLab Prop. ID 2012B-0001; PI: J. Frieman), which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation.
The DES data management system is supported by the National Science Foundation under Grant Numbers AST-1138766 and AST-1536171. The DES participants from Spanish institutions are partially supported by MICINN under grants ESP2017-89838, PGC2018-094773, PGC2018-102021, SEV-2016-0588, SEV-2016-0597, and MDM-2015-0509, some of which include ERDF funds from the European Union. IFAE is partially funded by the CERCA program of the Generalitat de Catalunya. Research leading to these results has received funding from the European Research Council under the European Union’s Seventh Framework Program (FP7/2007-2013) including ERC grant agreements 240672, 291329, and 306478. We acknowledge support from the Brazilian Instituto Nacional de Ciênciae Tecnologia (INCT) do e-Universo (CNPq grant 465376/2014-2).
This manuscript has been authored by Fermi Research Alliance, LLC under Contract No. DE-AC02-07CH11359 with the U.S. Department of Energy, Office of Science, Office of High Energy Physics.
@@ Appendix A Verification Tests
In Tables 4 and 5, we present the posterior distributions when we divide our sample into two subsets based on the redMaPPer central galaxy probability (). We use two samples with (Table 4) and (Table 5). When comparing these posterior distributions to those presented in Table 3, we see that all measurements are consistent within 1. This provides additional evidence that our results are not biased by miscentring or incorrect BCG identification, as BCGs with low are more likely to be miscentred or misidentified as the BCG.

### v2

@@ 7 Conclusion
Going forward, there remain many research questions we aim to explore to improve our observational understanding and characterisation of the BCG+ICL. In particular, understanding how the BCG+ICL stellar mass and radial distribution evolves with redshift and varies with wavelength. In future analysis using the Euclid Mission (as demonstrated by Kluge et al. (2024)) and the Rubin Observatory’s LSST (as discussed in Montes (2019) and Brough et al. (2024)), we plan to extend this type of analysis out to higher redshifts () and into the IR regime to investigate the formation and evolution of the BCG+ICL system and how its correlation with halo mass evolves. Moreover, given the correlation between M14 and ICL fraction, we suggest that characterizing large M14 systems in Euclid and LSST, or targetting them with JWST, may provide information about the earliest appearances of the ICL. Additionally, we aim to determine whether the correlation between M14 and BCG stellar mass persists out to high redshift and whether this is intrinsic (i.e., is it in place prior to ex-situ growth) or results entirely from hierarchical formation.
@@ Acknowledgements
JBGM would like to thank the anonymous referee for all of their help and feedback, which significantly strengthened this manuscript. JBGM would like to thank Nina Hatch, the members of the NottICL research group, and the participants in the ISSI ICL working group for many useful discussions about the nature of ICL and measurement techniques. JBGM gratefully acknowledges support from the Leverhulme Trust.
@@
The Collaborating Institutions are Argonne National Laboratory, the University of California at Santa Cruz, the University of Cambridge, Centro de Investigaciones Energéticas, Medioambientales y Tecnológicas-Madrid, the University of Chicago, University College London, the DES-Brazil Consortium, the University of Edinburgh, the Eidgenössische Technische Hochschule (ETH) Zürich, Fermi National Accelerator Laboratory, the University of Illinois at Urbana-Champaign, the Institut de Ciències de l’Espai (IEEC/CSIC), the Institut de Física d’Altes Energies, Lawrence Berkeley National Laboratory, the Ludwig-Maximilians Universität München and the associated Excellence Cluster Universe, the University of Michigan, NSF NOIRLab, the University of Nottingham, The Ohio State University, the University of Pennsylvania, the University of Portsmouth, SLAC National Accelerator Laboratory, Stanford University, the University of Sussex, Texas A&M University, and the OzDES Membership Consortium.
Based in part on observations at NSF Cerro Tololo Inter-American Observatory at NSF NOIRLab (NOIRLab Prop. ID 2012B-0001; PI: J. Frieman), which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation.
The DES data management system is supported by the National Science Foundation under Grant Numbers AST-1138766 and AST-1536171. The DES participants from Spanish institutions are partially supported by MICINN under grants PID2021-123012, PID2021-128989 PID2022-141079, SEV-2016-0588, CEX2020-001058-M and CEX2020-001007-S, some of which include ERDF funds from the European Union. IFAE is partially funded by the CERCA program of the Generalitat de Catalunya.
We acknowledge support from the Brazilian Instituto Nacional de Ciênciae Tecnologia (INCT) do e-Universo (CNPq grant 465376/2014-2).
This document was prepared by the DES Collaboration using the resources of the Fermi National Accelerator Laboratory (Fermilab), a U.S. Department of Energy, Office of Science, Office of High Energy Physics HEP User Facility. Fermilab is managed by Fermi Forward Discovery Group, LLC, acting under Contract No. 89243024CSC000002.
@@ Appendix A Verification Tests
In Tables 5 and 6, we present the posterior distributions when we divide our sample into two subsets based on the redMaPPer central galaxy probability (). We use two samples with (Table 5) and (Table 6). When comparing these posterior distributions to those presented in Table 4, we see that all measurements are consistent within 1. This provides additional evidence that our results are not biased by miscentring or incorrect BCG identification, as BCGs with low are more likely to be miscentred or misidentified as the BCG.
@@
Appendix B Example Posterior
In the MCMC analysis described in Section 4, we note that our output is given by the 2D posterior distribution. Since we run a number of MCMC analyses in this work, we choose not to display each individual posterior distribution because they are quite similar and can be more concisely summarized in Tables 4, 5, 6, 7. However, here, we present an example posterior distribution for the BCG+ICL region with M14 included. Figure 8 highlights the covariance that exists between and , which is reduced by removing the median values, and also that the values presented in each table represent the median and 1 values from the posterior distribution.
Affiliations
1 School of Physics and Astronomy, University of Nottingham, Nottingham, NG7 2RD, UK
2 NSF’s National Optical-Infrared Astronomy Research Laboratory, 950 Cherry Ave., Tucson, Az 85719, USA
3 Observatório Nacional, Rua Gal. José Cristino 77, Rio de Janeiro, RJ - 20921-400, Brazil
4 Fermi National Accelerator Laboratory, P. O. Box 500, Batavia, IL 60510, USA
5 Hamburger Sternwarte, Universität Hamburg, Gojenbergsweg 112, 21029 Hamburg, Germany
6 Wits Centre for Astrophysics, School of Physics, University of Witwatersrand, Private Bag 3, 2050, Johannesburg, South Africa
7 Astrophysics Research Centre, School of Mathematics, Statistics, and Computer Science, University of KwaZulu-Natal, Westville Campus, Durban 4041, South Africa
8 Laboratório Interinstitucional de e-Astronomia - LIneA, Rua Gal. José Cristino 77, Rio de Janeiro, RJ - 20921-400, Brazil
9 Department of Physics, University of Michigan, Ann Arbor, MI 48109, USA
10 Institute of Cosmology and Gravitation, University of Portsmouth, Portsmouth, PO1 3FX, UK
11 Department of Physics & Astronomy, University College London, Gower Street, London, WC1E 6BT, UK
12 Instituto de Astrofisica de Canarias, E-38205 La Laguna, Tenerife, Spain
13 Institut de Física d’Altes Energies (IFAE), The Barcelona Institute of Science and Technology, Campus UAB, 08193 Bellaterra (Barcelona) Spain
14 Centre for Extragalactic Astronomy, Durham University, South Road, Durham, DH1 3LE, UK
15 Centro de Investigaciones Energéticas, Medioambientales y Tecnológicas (CIEMAT), Madrid, Spain
16 Department of Physics, IIT Hyderabad, Kandi, Telangana 502285, India
17 Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Dr., Pasadena, CA 91109, USA
18 Institute of Theoretical Astrophysics, University of Oslo. P.O. Box 1029 Blindern, NO-0315 Oslo, Norway
19 Kavli Institute for Cosmological Physics, University of Chicago, Chicago, IL 60637, USA
20 Instituto de Fisica Teorica UAM/CSIC, Universidad Autonoma de Madrid, 28049 Madrid, Spain
21 Department of Physics and Astronomy, University of Pennsylvania, Philadelphia, PA 19104, USA
22 University Observatory, Faculty of Physics, Ludwig-Maximilians-Universität, Scheinerstr. 1, 81679 Munich, Germany
23 Center for Astrophysical Surveys, National Center for Supercomputing Applications, 1205 West Clark St., Urbana, IL 61801, USA
24 Department of Astronomy, University of Illinois at Urbana-Champaign, 1002 W. Green Street, Urbana, IL 61801, USA
25 School of Mathematics and Physics, University of Queensland, Brisbane, QLD 4072, Australia
26 Santa Cruz Institute for Particle Physics, Santa Cruz, CA 95064, USA
27 Center for Cosmology and Astro-Particle Physics, The Ohio State University, Columbus, OH 43210, USA
28 Department of Physics, The Ohio State University, Columbus, OH 43210, USA
29 Center for Astrophysics Harvard & Smithsonian, 60 Garden Street, Cambridge, MA 02138, USA
30 Australian Astronomical Optics, Macquarie University, North Ryde, NSW 2113, Australia
31 Lowell Observatory, 1400 Mars Hill Rd, Flagstaff, AZ 86001, USA
32 LPSC Grenoble - 53, Avenue des Martyrs 38026 Grenoble, France
33 Institució Catalana de Recerca i Estudis Avançats, E-08010 Barcelona, Spain
34 Department of Physics, Carnegie Mellon University, Pittsburgh, Pennsylvania 15312, USA
35 Kavli Institute for Particle Astrophysics & Cosmology, P. O. Box 2450, Stanford University, Stanford, CA 94305, USA
36 SLAC National Accelerator Laboratory, Menlo Park, CA 94025, USA
37 Department of Physics, Northeastern University, Boston, MA 02115, USA
38 Physics Department, Lancaster University, Lancaster, LA1 4YB, UK
39 Computer Science and Mathematics Division, Oak Ridge National Laboratory, Oak Ridge, TN 37831
40 Argone National Laboratory, 9700 S. Cass Avenue, Lemont, IL 60439, USA
41 Cerro Tololo Inter-American Observatory, NSF’s National Optical-Infrared Astronomy Research Laboratory, Casilla 603, La Serena, Chile
42 Department of Astronomy, University of California, Berkeley, 501 Campbell Hall, Berkeley, CA 94720, USA
43 Lawrence Berkeley National Laboratory, 1 Cyclotron Road, Berkeley, CA 94720, USA
44 School of Physics and Astronomy, University of Southampton, Southampton, SO17 1BJ, UK